OpenStudy (anonymous):

Maria and Charlie can deliver 80 papers in 4 hours. How long would it take them to deliver 46 papers?

hero (hero):

Since the number of hours increase as the number of papers increase, then we know this is a direct proportion so using the direct proportion formula we have: Direct Proportion Formula \[\frac{y}{x} = k\] In the first case, we find the constant of proportionality y = Number of papers = 80 x = Number of hours = 4 k = constant of proportionality \[\frac{80}{4} = k\]\[20 = k\] In the second case, we find the number of hours, x y = 46 x = ? k = 20 We use the same formula but plug in the values for the second case: \[\frac{46}{x} = 20\]\[\frac{46}{20} = x\]\[2.3 = x\] Thus, 46 papers can be delivered in 2.3 hours.